Bay Company sold$1,00,000 of 8%,10-year bonds at 97 on January 1,2021.The bonds were dated January 1,2021 and pay interest on June 30 and December 31.If Bay company uses the
straight-line amortization,what would the total interest expense recognized for the bond issue over its full term?
Answer #1

Total amount of bonds ($100,000) x 8% x Twice a year payment 6 months / 12 months)

100000*8%*6/12 = 8000*6/12 = $4000 per year interest expense

Interest expense period = 10 years

Interest expense = 4000*10 = $40000

The bond pays interest every 6 months on June 30 and December 31. We will amortize the discount using the straight-line method meaning we will take the total amount of the discount and divide by the total number of interest payments. In this question the discount amortization will be $3000 discount amount / 20 interest payment (10 years x 2 interest payments each year). The entry to record the semi-annual interest payment and discount amortization would be:

3000/20 = $150*10 years = 1500

So, the total Interest expense is $40000+ $1500 = $41500